OpenStudy (anonymous):

Find the standard form of the equation of the parabola with its vertex at the origin and with point (3,6).

OpenStudy (whpalmer4):

Vertex form is \[y=a(x-h)^2+k\]with vertex at \(h,k\) and \(a\) a constant you set to make the parabola fit your point. Then rewrite in standard form which is \[y=ax^2+bx+c\] (Note that a,b,c are not necessarily the same, I'm just showing the general form!)

OpenStudy (anonymous):

I need it in standard form like x^2=4py :/

OpenStudy (whpalmer4):

So rearrange it into that form!

OpenStudy (anonymous):

There is a in one equation and p in another though

OpenStudy (whpalmer4):

It all turns into numbers to fit the parabola.

OpenStudy (whpalmer4):

Where is your vertex?

OpenStudy (anonymous):

Vertex is 0,0

OpenStudy (whpalmer4):

Okay, h=0,k=0, plug those into my first formula, what do you get?

OpenStudy (anonymous):

y=a(x−h)2+k 6 =a(3-0)^2+0 6= 9a a= 2/3

OpenStudy (whpalmer4):

So what is your final equation?

OpenStudy (anonymous):

y=2/3x^2

OpenStudy (whpalmer4):

And with x^2 all alone on one side, that is?

OpenStudy (anonymous):

x^2 = 3/2y OMG THANKS!! :)

OpenStudy (whpalmer4):

Best to write that as (2/3)x^2, btw, just so there is no doubt about the meaning

OpenStudy (whpalmer4):

Yeah, and my formula will work even if the vertex is elsewhere...try it and see :-)

OpenStudy (anonymous):

do you mind helping me with another question?

OpenStudy (whpalmer4):

Sorry, I have to leave. Post and tag me, and I'll have a look when I return. Hopefully someone will help you before that.